Porchy

Pacman Diagnostic ROM

 General  Comments Off on Pacman Diagnostic ROM
May 292020
 

Philmur has sent me his Pacman diagnostic ROM.

Pacman test ROM – philmurr 05/20

Fits at location 6E

Tests programme, colour, video and sprite RAM 1 (note sprite RAM 2 is write-only)
Displays sprites dropping from screen to show sprite select and addressing is working (tries to show all 8 but most boards only have 6)
Shows real-time status of inputs and DIPs

Colour RAM – 4P, 4L
Prog RAM – 4R, 4M
Video RAM – 4N, 4K

Sprite RAM 1 – part of 4R,4M but may make it clearer to people when testing

File can be downloaded in the “Downloads/ROM files/Diagnostic” menu

Thanks to Philmur for his hard work and for sharing this

 Posted by at 4:32 pm
Apr 132020
 

Yet another Operation Wolf repair and this one gave me trouble.
First test we had watchdogging. All ROM’s checked out fine in the programmer and seeing nothing else obvious I removed the work RAM both of which failed.

Now on boot up I got a screen seemingly running game but all the screen was garbage.

Probing the data lines of the screen RAM show pretty much all of them across two RAM chips were dead. I desoldered and replaced them both.

Now I got a running game but the colours were messed up

There are two 74LS245’s attached to the data lines of the colour RAM at location IC78 & IC90. Probing these revealed stuck bits on IC78. Replacing this gave me perfect results on the colour test screen in test mode but dull colours in game. Other issues in-game showed me all the sprites were missing and the screen doesn’t scroll when entering a game.

Stupidly I chose to ignore the colour issue and concentrate on the other two (I will get back to this later).
The sprite and scrolling issue worried me because these areas are related to the PC080SN and PC0900J custom chips.

Paulcan69 sent me a scrapper PCB so I could rob the customs off it if needed so I started out by replacing the PC0900J.

Replacing this brought my sprites back but the scrolling issue was still present.
I didn’t enjoy replacing that IC so I started checking around to see if the scrolling issue could lie somewhere else.
I found nothing wrong so removed the PC080SN custom and replaced it with the spare but it made absolutely no difference. A lot of work time and effort for nothing.
I spent time over the course of the next few days looking into this but found nothing.
Eventually I got sick of looking at the washed out colours on screen and replaced IC90 to fix the last of the colour issues.
I couldn’t really believe it when this fixed my scrolling issue as well.
I still don’t understand why this chip caused the scrolling issue but I’m really glad it did.
All issue are now fixed on this board.

Onto the sound PCB.
The original sound PCB that came with this board was a complete write off. Every chip I pulled from this board was broken and both amps were also burning hot.
Among all the other boards that Paulcan69 had sent me he also kindly sent me a spare sound PCB for Muddymusic to have.
There was a sticker on this board saying “No Sound”.
On first test I found that the music played but the volume was really low. Turning the volume up to the maximum made the sound audible but not nearly enough.
I checked the resistance of the volume pot and found that if you move the pot off maximum setting then it read megaohms. I tried cleaning it but it didn’t work so I replaced it.
This made the sounds much better but like with most of the others I’ve repaired some of the samples were scrambled and didn’t stop playing when they should.

I quickly found another 74LS688 that was dead. Replacing this made the sounds stop when they should but the samples on channel A were scrambled.
Looking at the schematics shows where the sample data goes.

Using an audio probe I checked the output at pin 10 of the 5205. The sound was the same. This ruled out issues caused by bad caps and things like that.
Next I checked the 574 at IC33. Both the input and output looked good so moved onto the 74LS157 at IC 44.
Probing all the outputs with a logic probe showed pin 9 was stuck HIGH

Replacing this fixed all the samples.
That’s the last one fixed
Big thanks to Paulcan69 for the spares board. Without it this board would be scrap too.

 Posted by at 9:54 am

Operation Wolf repair log #4

 PCB Repair Logs, Repair Logs  Comments Off on Operation Wolf repair log #4
Apr 112020
 

Another Operation Wolf repair!
This one belongs to Frothmeister on UKVAC.

Again, we have a nice little fault label

So to start with there was no sound. I did a couple of signal checks to see if commands were being written to various IC’s but saw none of them so I just removed the RAM which of course failed.

Fitting a new RAM brought the sounds back but the samples were garbage

Samples are all stored in the 40 pin MASKROM.

I attempted to read the MASKROM as a 27C400 EPROM in my programmer but it gave me errors

As I had a scrap sound board here I swapped this ROM out and retested.
Samples were now restored but I had the same issue as in a previous repair where the sounds started playing and never stopped.

Time to check out the comparators.
I quickly found the comparator with a stuck output at location IC18

Replacing this fixed my faulty sample playback for this one but I found another by playing sample 2E.
Using the same technique I found another dead 688 at IC39.

To determine which channel the game uses for various samples I use the wonderful MAME debugger.
If you look at the memory map for the sound CPU in the MAME source you can see the addresses that get written to to set the START and STOP values.

Address ranges $B000 – $B006, $C000 – $C006, $D000 – $D006 & $E000 – $E006 are what we are concerned with so set some breakpoints in the debugger and fire up the Operation Wolf test menu (dont forget to set the debugger focus onto the sound CPU).
Play the sounds that cause issues and see which address gets written to.

Now look at the schematics

From here we need to work out with signal is address $C000.
A15 & A14 give us an address of $C000. the 74LS138 at location IC14 gives us the enable line for the channel. A15 enables the 138 itself and A14 selects the relevant output, in this case its input C which enables (active LOW) pin 11. This in turn enables IC54 which as you can see from the signal names is channel B.
Following the schematics further we find the 74LS688 comparators responsible for this

Probing the only output on those comparators (pin 19) and playing the sample shows me which one isn’t toggling. In this case it was IC39.

All fixed

 Posted by at 3:47 pm

Operation Wolf repair log #3

 PCB Repair Logs, Repair Logs  Comments Off on Operation Wolf repair log #3
Apr 102020
 

Another few Operation Wolf boards in at the moment. This one is from Unit504.

There is a nice little tag on this telling me the faults which really helps me keep all these boards together and where to start looking.

First job was to look into the gun shot register issue.
I did a quick test and could see the screen flash when the trigger was pulled so I knew that wasnt the fault
Following the circuit I come to the 74F74 IC which is used to latch the gun co-ordinates.

I normally wouldn’t start looking at this part but it was already socketed so decided to pull the chip and test it. It failed an out of circuit test.

I replaced this with a 74ALS74 and it seems to be be fine.
Now all the gun shots register as expected.

Next onto the sound.
None of the sounds were working at all and I didnt believe that all the seperate circuits for making sounds would be dead so looked a bit closer at the CPU side of things.
First off I checked the ROM and it dumped out fine.
Next the RAM. An inspection of the RAM showed signs of corrosion

I removed and tested it and thankfully it failed

Looks like that corroded pin has broken contact somewhere. Anyway, replacing the RAM brought all the sound back to life and completed this repair.

Operation Wolf repair log #2

 PCB Repair Logs, Repair Logs  Comments Off on Operation Wolf repair log #2
Mar 272020
 

Got this board set from Paulcan69 to have a look at.
The faults reported were gun shots do not register & there was no sound.
I was told before hand that the gun issue seems to be a very common problem so I focused my attention on that first.
The input from the gun comes in from the sound PCB and passed straight down to the main PCB.
First I checked the signal from the gun was actually reaching where it should do

I confirmed that on a trigger press in game that the signal reaches pin 3 of the 74F74 at IC26 and also that pin 5 output latched.
Using the scope I could also confirm the operation of the 74LS75 at IC38.

The current XY screen draw position is tracked by a set of 74LS161 counters and fed into 74LS373 transparent latches.

When the trigger is pressed the the 74LS373’s are latched and the current XY values are sent to 3 x 74LS257 data selectors.

These selectors are read by the main CPU from address $3A0000 – $3A0001.
Using the scope syncing on pin 15 of the 74LS257’s I confirmed that some of the outputs on these were always logic HIGH.
I ended up replacing IC67 & IC83 which fixed the gun problem.

Next was the sound issue.
This one had no sound at all.
All the signals looked OK from the CPU side but nothing was ever being asked to play a sound.
Given the fact that the board uses the Sony brand RAM I just went right ahead and removed it as they have been very unreliable in recent times.
It failed in an out of circuit test using my programmer

Now the music was back but the samples did not.
I already had a pretty good grasp on how the sound worked from my previous repair so started checking the same areas as before but didn’t really find anything but then I spotted this

This is part of the feedback circuit for the op-amp. Replacing this gave me some samples back but they cut out before they finished playing and some didn’t play at all

I knew exactly where to look for this and quickly found the 74LS688 at location IC41 with its output stuck LOW.

I removed and tested which obviously failed

I initially thought I would have to put this repair on hold as I needed to order some 74LS688’s but turned out I had to wait less than 24hrs for Farnell to deliver.
With the new chip fitted I had my sounds playing for the right duration but all the samples on the A channel were still missing altogether.
Using the scope I traced the sound to the output of the M5M5205 ADPCM Speech Synthesis LSI at IC50 (schematics incorrectly refer to it as IC51)

I could see the signal going into the op-amp but nothing was coming out so I replaced it.

I now had an output from the op-amp but still no sound. Tracing this through on the PCB I came across another damaged capacitor (no idea how I missed this one as well)

Replacing this brought all the sound back and that’s another one complete.

 Posted by at 8:01 pm